geerky42 (geerky42):

I have a trouble understanding the representations of objects... Saying we have \[\cal{R}:\mathbb{Q}^+ \longrightarrow~~\mathbb{N}\times\mathbb{N}\]defined by \(\cal{R}\left(\dfrac{a}{b}\right) = (a, b)\), where \(a,b\in\mathbb{N}\) and \(b\neq0\) Is \(\cal{R} \) a function or not?

geerky42 (geerky42):

What so hard for me to understand is that we know that \(\dfrac{1}{2} = \dfrac{2}{4}\), so we could say that \(\cal{R}\) is not function because \(\cal{R} \left( \dfrac{1}{2}\right) = (1,2)\) and \((2,4)\). But \(\dfrac{2}{4}\) doesn't necessary exist in \(\mathbb{Q}^+\), since it already has \(\dfrac{1}{2}\), because, well, \(\dfrac{2}{4}\) is simplified to \(\dfrac{1}{2}\) right? What I said is probably sounds dumb, but it is ambiguous to me...

OpenStudy (zarkon):

it is not well defined because of the reason you stated 1/2 and 2/4 give different outputs Also, 2/4 is in \(\mathbb{Q}^+\)